# On the Intrinsic Self-Correction Capability of LLMs: Uncertainty and Latent Concept

### v1

@@ 2.3 The Collaborative Effect of Uncertainty and Latent Concept
In section 2.1 and 2.2, we examine how model uncertainty and the activated concept evolve as the self-correction process progresses towards convergence and improved performance. In this section, we empirically validate the collaboration between model uncertainty and activated concept in terms of driving LLMs towards increasingly better performance and converging eventually.
We utilize the text detoxification task and intervene in the self-correction process using negative instructions or sub-optimal self-correction instructions. Positive instructions are defined as those that guide LLMs towards the desired objective, such as producing less toxic outputs. Sub-optimal instructions, while characterized by greater uncertainty and a less robust activated concept compared to positive instructions, still steer the LLMs toward the desired objective. Conversely, negative instructions are those that counteract the objective, such as increasing the toxicity of the generated text. Details about the used instructions are available in Appendix D.7.
Figure 4: The uncertainty (left), toxic similarity (middle, the cosine similarity between a given hidden state and the toxic probing vector), and the toxicity score (right) by using positive, negative, and sub-optimal instructions. We report the performance by interacting with LLMs for 8 rounds. All sub-figures show a tendency towards convergence, including the results acquired with negative instructions.
Figure 4 illustrates how the uncertainty, toxic similarity (the cosine similarity between a given hidden state and the toxic probing vector), and toxicity score change across 8 rounds, by using different instructions. The following key observations can be found: Obs.1. Regarding the toxicity score depicted in the right sub-figure, both positive and sub-optimal instructions lead LLMs to generate less toxic responses, with the positive instructions demonstrating superior performance. Conversely, negative instructions result in increasingly toxic responses. Obs.2. In terms of uncertainty, all types of instructions contribute to a reduction in uncertainty, with a tendency towards convergence observed by the final round. This suggests that the decrease in uncertainty is independent of the nature of the given instruction. Obs.3. The toxic similarity (middle sub-figure) quantifies the extent to which the activated concept is toxic. The positive instruction activates the least toxic concept, followed by the sub-optimal instruction, while the negative instruction significantly increases the toxicity of the response. Obs.4. The correlation between the toxic level of the activated concept and the nature of instruction is significant. These observations demonstrate that uncertainty and the activated concept are two influential factors for the self-correction performance.
Figure 5: The uncertainty (left), toxic similarity (middle, the cosine similarity between the hidden state and the toxic probing vector), and toxicity score (right) by intervening in the self-correction progress in the 2nd, 5th, and 8th rounds. The Origin represents the positive-instructed self-correction process without intervention of negative instructions. We report the metric gap of the current round to previous round. The intervention of negative instructions introduces perturbation with all considered metrics, and the change of toxicity score is highly correlated to that of uncertainty and toxic similarity.
To achieve a more fine-grain analysis, we intervene self-correction process with negative instructions in the $2^{nd},5^{th}$ and $8^{th}$ round. Figure 5 presents what is the difference of the current round to previous round if intervention with negative instructions happens. The toxicity score shown in the right subfigure shows that intervene with negative instruction can make the responses more toxic since the toxicty score of previous round is smaller than that of the current round. This is also reflected in the toxic similarity sub-figure (middle), the negative instruction makes the activated concept be more toxic than the concept in the previous round. This demonstrates the significant impact of instruction types on the toxicity levels of the activated concepts.
The reported uncertainty in the left sub-figure shows similar evidence for the intervention. In the $2^{nd}$ round, the introduction of negative instruction causes a smaller decline of uncertainty compared to that without the negative instruction. For the $5^{th}$ and $8^{th}$ round, the original self-correction process has already been close to convergence without too many fluctuations, however, the uncertainty of the intervention process decreases with a large magnitude in the round of 5. Those empirical evidence advocates the association between the perturbation of self-correction performance and that of both uncertainty and the activated concept.
In summary, uncertainty and activated positive concept are critical variables in the self-correction process, and our intervention experiments indicate that they are fundamentally correlated with self-correction performance.
@@ 2.4 Theoretical Analysis Towards the Convergence of Self-correction
Section 2.1 and 2.2 show empirical evidence about the collaborative effects of uncertainty and activated concept to the convergence of self-correction. In this section, we present a straightforward yet inspiring mathematical formulation of self-correction within the framework of multi-round question-answering. Our formulation explicitly illustrates the significant role of activated concepts to the convergence of self-correction and uncertainty, and provides insights into enhancing self-correction performance by devising effective pre-hoc self-correction instructions for the first interaction round.
Let the input question be denoted as $x$, an individual instruction as $i\in\mathcal{I}$ wherein $\mathcal{I}$ represents the set of all possible self-correction instructions that can yield the desired and harmless responses given a task. Let $y$ denotes the output of LLMs. For the $t^{th}$ round of interaction, the input sequence to LLMs $f$, parameterized with $\theta$, is represented as $q_{t}=$ ($x,i_{0},y_{0},i_{1},y_{1},i_{2},y_{2},\dots,i_{t}$) for $t>2$ and the response $y_{t}=f_{\theta}(q_{t})$.
[…]
@@
Given the assumption that the measurement over the response depends on the activated concept of the associated inputs to LLMs. The objective of self-correction can be interpreted as:
$p(\mathcal{M}(y_{t}|y_{t-1})\geq\mathcal{M}(y_{t-1}|y_{t-2}))\Rightarrow\sum_{k=0}^{t}p(C^{*}|q_{k})\geq\sum_{k=0}^{t-1}p(C^{*}|q_{k-1})$ (1)
This is always held once $p(C^{*}|q_{k})>0$, the equal sign stands for the convergence of self-correction performance. By delving into each term of probability we can have:
$\begin{split}p(C^{*}|q_{0})=p(C^{*}|x)p(C^{*}|i_{0})&=\mathrm{c}_{x}\mathrm{c}_{i},k=0\\ p(C^{*}|q_{1})=p(C^{*}|x)p(C^{*}|i_{0})p(C^{*}|y_{0})p(C^{*}|i_{1})&=\mathrm{c}_{x}\mathrm{c}_{i}\mathrm{c}_{y}\mathrm{c}_{i},k=1\\ \vdots\\ p(C^{*}|q_{k})=p(C^{*}|x)p(C^{*}|i_{0})p(C^{*}|y_{0})\dots p(C^{*}|i_{k})&=\mathrm{c}_{x}\underbrace{\mathrm{c}_{i}\mathrm{c}_{y}\mathrm{c}_{i}\mathrm{c}_{y}\dots\mathrm{c}_{i}\mathrm{c}_{y}}_{{(\mathrm{c}_{i}}{\mathrm{c}_{y}})^{t-1}}\mathrm{c}_{i},k=t(t>1)\\ \end{split}$ (2)
Since all probability constant $\mathrm{c}$ is positive and is less than 1, we can have $\lim\limits_{t\to\infty}p(C^{*}|q_{t})=0$, and $p(C^{*}|q_{k})=(\mathrm{c}_{i}\mathrm{c}_{y})^{t-1}p(C^{*}|q_{0})<p(C^{*}|q_{0})$. Therefore the positive effects of self-correction instructions degrades as the interaction round increases, and, in practical scenarios, we observe the probability of positive concept $C^{*}$ drops fast and approaches 0 after only several rounds. Our formulation indicates (a) the dominant gain of the self-correction instruction in the first round, a.k.a. pre-hoc instruction. (b) the significant difficulty of figuring out a feasible self-correction instruction, the instruction should compensate the negativeness of noise introduced by other information, e.g. input question $x$, generated response $y$. Such theoretical analysis can also verify the empirical observations in [20, 39], advocating the importance of proper instruction in the first round when taking the computational budget of self-correction into consideration.
However, we can not overlook the possibility that $x,y,i$ can also activate negative concepts, potentially leading to undesired and harmful outcomes, thereby invalidating our assumptions in the mathematical formulation presented above. Our empirical observations in Figure 3 and 4 shows the decline of activated concept in the second round for most tasks, but the decline is slight and does not hurt the self-correction performance seriously.
[…]
The identical equation is hold since $\theta$ is fixed in practice. Equation 3 indicates the convergence of concept helps the uncertainty converge as well, implying the interaction between those two variables from a theoretical view. For cases where $p(C^{*}|q)<p(\hat{C}|q)$, though the negative concept is activated but the certainty is not decreased, explaining why the injected negative instruction at the round of 2, 5 or 8 does hinder the decrease of uncertainty (Figure 5).
In summary, the convergence of activated positive concepts drives the convergence of uncertainty and self-correction, and the first-round instruction is the most crucial for the final performance.
3 Practical Applications Inspired by Analytical Insights
Self-correction is a test-time technique that utilizes instructions to explicitly guide LLMs towards desired behaviors. In Section 2, our analytical insights indicate that decreasing model uncertainty and activating positive concepts are two crucial factors jointly contributing to success. We hypothesize that the importance of these two factors can be generalized to other applications of LLMs.
To demonstrate the generalizability of these insights, we investigate whether fine-tuning, in addition to self-correction, can also benefit from our findings related to model uncertainty and activated positive concept. To this end, we propose a data selection method for task-agnostic gender bias mitigation, implemented through counterfactual data augmentation [40], which involves the following key steps: (1) activated positive concept: filtering out sentences without any gendered words and flipping gendered words in the debiasing corpus, (2) model uncertainty: calculating the uncertainty of a GPT2-XL [^3] https://huggingface.co/openai-community/gpt2-xl model [36] for each sample in the debiasing corpus, and (3) categorizing the debiasing samples into two groups based on their uncertainty and fine-tuning a GPT2-XL model with each group. To estimate the uncertainty of LLMs regarding a debiasing sample [36], we incorporate each sample into a question-answer task. Following the methodology of [41], we instruct the LLMs to determine whether the sample statement is biased or not. Further details regarding the experimental setting are provided in Appendix D.5. Notably, this data selection method can be easily adapted for various fine-tuning or instructing purposes for LLMs. We utilize task-agnostic debiasing as a concrete example to illustrate the effectiveness of this approach.
Figure 6: Social bias mitigation for gender stereotypes in GPT2-XL. The language model modeling score (the likelihood of a natural language sequence generated by the given LLMs, the higher the better) and the StereoSet score are reported (the score is shown by the difference to the absolute fair of 0.5, the lower the better). The performance of fine-tuning progress associated with dataset of certainty is shown in blue, and the performance of uncertainty dataset is colored in orange.
Figure 6 presents the performance of the language modeling score (the likelihood of a natural language sequence generated by the given LLMs) and the approximated StereoSet score (the difference between the real StereoSet score and the absolute fairness of 0.5), during the fine-tuning process. Samples from the certainty group exhibit less uncertainty compared to those from the uncertainty group, although the evaluated LLM are not entirely certain about the certainty group. Regarding the language modeling score, the curve for the certainty group consistently remains above that of the uncertainty group across all fine-tuning epochs, indicating that the LLMs are more adept at fitting data they are more certain about. Overall, the StereoSet score for the certainty group is lower than that for the uncertainty group, with the gap widening as epochs progress, further proving that higher uncertainty prevent LLMs from an optimal fitting.
Previous studies have shown that the nature of this uncertainty-driven data selection approach is the selected samples are more informative [42, 43]. With the empirical evidence, we argue that this should be a principle for LLMs as well, and we believe this can used for understanding why self-instruct [30] and self-alignment [31] can outperform human-annotated datasets. The generated samples from LLMs themselves for self-instruct and self-alignment purpose are from the approximated distribution of LLMs’ knowledge of pre-training corpora. Therefore, those samples enjoy lower uncertainty than those annotated by humans. We leave this for future research.
@@ 4 Related Work
Self-correction is the capability of LLMs that allows them to modify their outputs based on instructions or external feedback. Such ability enables LLMs to adjust their responses for improved accuracy, relevance, and coherence, helping LLMs more effective in various applications. Proper-designed self-correction instruction has revealed empirical success in various application scenarios, e.g., machine translation [16], code generation [15], social bias mitigation [41]. Self-correction techniques [44] can be roughly categorized into (1) instruction-based, utilizing vanilla natural language instruction and intrinsic self-correction capability of the LLM (2) external-feedback based one, relying on an external verifier to provide external feedback. Our paper focuses on the intrinsic capability of LLM and the instruction-based self-correction techniques while leaving the external ones as important future work. More discussions on related work can be found in Appendix A.
5 Conclusion & Future directions
This work provides crucial insights into the effectiveness of self-correction across multiple tasks. The key reason for the success is that the proper-designed instruction can activate the correct concept with increasing certainty. Inspired by the importance of the concept and certainty, we introduce a simple but effective method to select fine-tuning data for social bias mitigation, demonstrating positive social impact. More discussion on the broader impact can be found in Appendix C.
Our paper focuses on the intrinsic self-correction capability of LLMs while leaving the self-correction with external feedback, e.g., tools, human preference, and retrieval knowledge, as the future work. For instance, [45] points out that the LLM must hallucinate on certain types of fact knowledge, only relying on its intrinsic capability. Retrieval-Augmented Generation (RAG) technique, which retrieves relevant information from an external knowledge base to ease the LLM generation, is essential to alleviate the hallucination phenomenon. Beyond the intrinsic capability, how the LLMs adaptively take advantage of the external feedback [46] can be an interesting future direction.
Our paper leaves the reasoning task requiring to generate intermediate explanation step as a future direction. A debate is currently remained on the effectiveness of self-correction techniques on reasoning tasks. [20] reveals that LLMs struggle to correct their responses while [21] observes consistent improvement. According to our new insights, we conjecture that the current evaluation which only examines the final performance may not be optimal to access the effectiveness of the self-correction technique. [47] observes that the intermediate explanation step may not faithfully reflect the true reason process for the LLM predictions. Consequently, the correctness of intermediate steps may not be relevant to the correctness of the final prediction. The self-correction may help to correct the intermediate step but hurt the final prediction. Therefore, an effective self-correction process with correct intermediate steps may not lead to performance improvement. There remains mystery on the intersection between the self-correction and reasoning capabilities of LLM. Nonetheless, we currently do not have sufficient GPU resources to analyze larger LLMs with proper reasoning capability.
In this paper, we only explore how the activated concept and uncertainty interplay for the convergence of self-correction. However, the causality between those two driving forces is not clear. Further research could figure out the causality and come up with solutions to have better self-correction instructions. The activated concept is not formally defined and is far from practical usage of this feature for better self-correction performance. It might be the consequence of activated subnetworks [48], how to efficiently detect subnetworks and associate them to concepts would be an interesting research question.
Appendix A Related work
A.1 Uncertainty estimation
@@
A.2 More discussion on Self-correction
Moreover, our paper shows correlation with [20], a recent empirical analysis paper on the self-correction technique. Our paper can provide additional explanation on phenomenons found in [20]. [20] finds that LLMs struggle to amend their prior responses where the GPT3.5 0301 version almost always believes its initial response is correct. We hypothesize such phenomenon is due to the model initial response reach a high certainty with no further modification in the later stage. [20] also finds that enhancement attributed to self-correction in certain tasks may stem from an ill-crafted initial instruction that is overshadowed by a carefully-crafted feedback prompt. Our theoretical analysis in Section 2.4 further explain the effectiveness of the initial prompt.
A.3 Instruction following
The self-correction technique is a well-known instruction-based method that requires LLMs to have a strong capability to follow instructions. However, vanilla LLMs may not be good at following instructions from humans [51]. To address this issue, recent LLMs have been equipped with instruction tuning techniques [52, 11, 51], which utilize templates and response pairs in text-to-text format [53] and show effectiveness on following instruction to unseen tasks. More recently, advanced instruction tuning techniques [54, 55, 56] have been developed to acquire labor-free, task-balancing, and large-scale instruction-following data. To quantify the instruction following capability, [57, 58] collect datasets towards scalable and cost-effective evaluation methods. To quantify instruction-following capability, datasets for scalable and cost-effective evaluation methods have been conducted [59, 60, 61], which evaluates on adversarial, counterfactual, and unnatural instruction following scenarios. Our paper focuses on how to better utilize the existing instruction following capability on self-correction tasks.
Appendix B Additional Experimental Results
B.1 Visualization Results

### v2

@@ 6 The Essential Force for Convergence
In Sections 4 and 5, we examined how model uncertainty and the activated concept evolve as the self-correction process progresses towards convergence and improved performance. In this section, we empirically and theoretically validate the collaboration between model uncertainty and activated concept in terms of driving LLMs towards increasingly better performance and eventual convergence.
In Section 6.1, we present empirical evidence establishing a dependent link between latent concepts and model uncertainty through a simulation task, wherein we utilize concept-relevant signals to predict changes in model uncertainty. Based on this dependence relationship, in Section 6.2, we provide a mathematical formulation demonstrating how self-correction instructions guide model uncertainty toward improved calibration, ultimately leading to more stable and converged performance.
6.1 The Dependence between Concept and Model Uncertainty
Referring to Equation 1, we present the mathematical formulation that links concepts to model uncertainty, specifically $p(c|q_{t},\theta)$. However, another term, $p(y_{t}|c,q_{t},\theta)$, also contributes to the overall uncertainty. To empirically validate the strong causal relationship between concept and uncertainty, we propose a simulation task framed as a binary classification problem. This task leverages the concept shift across any two self-correction rounds to predict whether uncertainty will increase or decrease.
Task Description. For each self-correction trajectory, we randomly sample two rounds of interaction and get the concepts ($c_{1},c_{2}$) and uncertainty values ($u_{1},u_{2}$). Please note the concept is represented as the cosine distance between each layer-wise hidden state and the probing vector, so $c_{1}\in\mathbb{R}^{l}$ and $c_{2}\in\mathbb{R}^{l}$, where l is the number of transformer layers. $u_{1},u_{2}$ are acquired through the semantic uncertainty [34] as introduced in Section 4. We leverage $u_{2}-u_{1}$ as the change of concept and the label is set as 1 if $u_{2}-u_{1}$ is no larger than 0, otherwise the label should be 0.
In our implementation, we randomly sample 2,000 questions from the RealToxicityPrompts benchmark, using 1,600 for the training set and the remaining 400 for the test set. We employ a linear classification model (logistic regression [^6] https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.LogisticRegression.html) and conduct the experiment five times [^7] The seed set includes 1, 25, 42, 100, and 1000.. The model achieves an average accuracy of 83.18%, with a variance of 0.00024.
Given Equation 1 and the experimental results of the simulation task, we can conclude that there is a strong dependence between the activated concept and model uncertainty. In other words, the concept activated through self-correction instructions is a strong driving force for the change in model uncertainty.
@@ 6.2 Theoretical Analysis Towards the Convergence of Self-correction
Previous sections have shown empirical evidence about the model uncertainty, how concepts activate and evolve per the self-correction process, and how model uncertainty is dependent on the concept. In this section, we present a straightforward yet inspiring mathematical formulation of self-correction, to further reveal how instructions help performance converge from a theoretical point of view.
In the context of QA interaction, the goal of self-correction is to ensure that $\mathcal{M}(y_{t}|y_{t-1})\geq\mathcal{M}(y_{t-1}|y_{t-2})$ where $\mathcal{M}$ is a metric measuring some properties of a given output, such as non-toxicity, harmlessness. $y_{t}|y_{t-1}$. We have the independent assumption over question $x$, instruction $i$ and output $y$, e.g., $p(x,i,y)=p(x)p(i)p(y)$, and denote $p(C_{p}|x)=\mathrm{c}_{x}(0<\mathrm{c}_{x}<1)$ , $p(C_{p}|y)=\mathrm{c}_{y}(0<\mathrm{c}_{y}<1)$, $p(C_{p}|i)=\mathrm{c}_{i}(0<\mathrm{c}_{i}<1)$, $p(C_{p})=\mathrm{c}_{p}(0<\mathrm{c}_{i}<1)$, and $\mathrm{c}_{x}\ll\mathrm{c}_{i}$.
@@
Given the assumption that the measurement over the response depends on the activated concept of the inputs to LLMs. The objective of self-correction can be interpreted as:
$\mathcal{M}(y_{t})\geq\mathcal{M}(y_{t-1})\Rightarrow p(C_{p}|q_{k})>p(C_{n}|q_{k})\geq 0,~{}\forall k:0<k\leq t$ (2)
The equal sign stands for the convergence of self-correction performance, implying the self-correction performance would be stable since round $k$. Our empirical analysis in Section 5 provides evidence that the activated concept is the positive one $C_{p}$ as long as the injected instruction $i_{k}$ is relevant to the desired goal, i.e., less toxic, no gender bias. Therefore $p(C_{p}|q_{k})>p(C_{n}|q_{k})$ holds for any $k$. By delving into each term of probability we show how the activated concept changes as the interaction round progresses from 0 to $t$:
$\begin{split}p(C_{p}|q_{0})=\frac{p(C_{p}|x)p(C_{p}|i_{0})}{p(C_{p})}&=\frac{\mathrm{c}_{x}\mathrm{c}_{i}}{\mathrm{c}_{p}},k=0\\ p(C_{p}|q_{1})=\frac{p(C_{p}|x)p(C_{p}|i_{0})p(C_{p}|y_{0})p(C_{p}|i_{1})}{p(C_{p})}&=\frac{\mathrm{c}_{x}\mathrm{c}_{i}\mathrm{c}_{y}\mathrm{c}_{i}}{\mathrm{c}_{p}},k=1\\ p(C_{p}|q_{k})=\frac{p(C_{p}|x)p(C_{p}|i_{0})p(C_{p}|y_{0})\dots p(C_{p}|i_{k})}{p(C_{p})}&=\frac{\mathrm{c}_{x}\overbrace{\mathrm{c}_{i}\mathrm{c}_{y}\mathrm{c}_{i}\mathrm{c}_{y}\dots\mathrm{c}_{i}\mathrm{c}_{y}}^{{(\mathrm{c}_{i}}{\mathrm{c}_{y}})^{t-1}}\mathrm{c}_{i}}{\mathrm{c}_{p}},k=t(t>1)\\ \end{split}$ (3)
Since $\mathrm{c}_{p}$ is a constant, we can have $p(C_{p}|q_{k})=(\mathrm{c}_{i}\mathrm{c}_{y})^{t-1}p(C_{p}|q_{0})<p(C_{p}|q_{0})$. This implies that the effect of the positive concept activated by self-correction instructions degrades as the interaction round progresses. The overall effects of positive concepts converges at a typical round because, since this round, the probability $p(C_{p}|q_{k})\approx 0$ but $p(C_{p}|q_{k})>p(C_{n}|q_{k})$ which is guaranteed according to our empirical evidence about the irreversability property of activated concepts. This formulation explains why model uncertainty evolves towards convergence as shown in Figure 4.
In practical scenarios, we observe the performance of self-correction does not improve after only several rounds. Our formulation further demonstrates the substantial impact of the self-correction instruction in the first round, consistent with previous studies that highlight the importance of providing appropriate instructions in the first round [19, 47].
In conclusion, Equation 1 establishes the connection between the activated concept and model uncertainty, while Section 6.1 provides empirical evidence supporting the dependence between these two variables. We can therefore conclude that the converged uncertainty reported in Section 4 is driven by the convergence of activated positive concepts. This finding bridges the relationships among self-correction instructions, activated concepts, model uncertainty, calibration error, and the converged performance, as illustrated in the logical framework (Figure 2).
7 Discussions
[29] empirically demonstrates that intrinsic moral self-correction is superficial, as it does not significantly alter immorality in hidden states. Our study addresses the question of why intrinsic self-correction is still effective despite its superficiality. We exclude reasoning tasks from our analysis due to ongoing debates surrounding the effectiveness of self-correction in reasoning [19]. Intrinsic moral self-correction is a practical instance of the Three Laws of Robotics [48]; with this principle we expect AI can follow our abstract orders and take harmless actions. In this paper, we implement in-depth analysis in the context of toxic speech. This is partially because the toxicity can be directly inferred from languages and it is more straightforward to humans than other moral dimensions such as social stereotypes [46]. On the other hand, for toxic speech, we can leverage more tools for interpreting black-box models to understand intrinsic self-correction. Our research functions as a prototype to analyze the self-correction capability in other scenarios such as language agents [49, 50]. Among those applications of language agents, our analysis framework can also be applied by defining the concept as the intent or actions towards the goal of a specific agent.
@@ 8 Related Work
Self-correction is the capability of LLMs that allows them to modify their outputs based on instructions or external feedback. Such ability enables LLMs to adjust their responses for improved accuracy, relevance, and coherence, helping LLMs more effective in various applications. Proper-designed self-correction instruction has revealed empirical success in various application scenarios, e.g., machine translation [16], code generation [15], social bias mitigation [51]. Self-correction techniques [52] can be roughly categorized into (1) instruction-based, utilizing vanilla natural language instruction and intrinsic self-correction capability of the LLM (2) external-feedback based one, relying on an external verifier to provide external feedback. Our paper focuses on the intrinsic capability of LLM and the instruction-based self-correction techniques while leaving the external ones as important future work. Moreover, our paper shows correlation with [19], a recent empirical analysis paper on the self-correction technique. Our paper can provide additional explanation on phenomenons found in [19], which shows that LLMs struggle to amend their prior responses where the GPT3.5 almost always believes its initial response is correct. We hypothesize such phenomenon is due to the model initial response reach a high certainty with no further modification in the later stage. [19] also finds that enhancement attributed to self-correction in certain tasks may stem from an ill-crafted initial instruction that is overshadowed by a carefully-crafted feedback prompt. Our theoretical analysis in Section 6.2 further explain the effectiveness of the initial prompt.
@@
The instruction-following capability of LLMs is the foundation for self-correction. However, vanilla LLMs may not be good at following instructions from humans [54]. To address this issue, recent LLMs have been equipped with instruction tuning techniques [55, 11, 54], which utilize templates and response pairs in text-to-text format [56] and show effectiveness on following instruction to unseen tasks. More recently, advanced instruction tuning techniques [57, 58, 59] have been developed to acquire labor-free, task-balancing, and large-scale instruction-following data. To quantify the instruction following capability, [60, 61] collect datasets towards scalable and cost-effective evaluation methods. To quantify instruction-following capability, datasets for scalable and cost-effective evaluation methods have been conducted [62, 63, 64], which evaluates on adverserial, counterfactual, and unnatural instruction following scenarios.
9 Conclusion & Future Work
Conclusion. In this paper, we validate the convergence phenomenon of intrinsic self-correction across various tasks and LLMs/VLMs, and reveal that the effectiveness of intrinsic self-correction stems from reduced model uncertainty. Specifically, we show empirical evidence and theoretical formulation that the convergence of activated concepts by self-correction instructions drives the model uncertainty towards convergence, therefore motivating LLMs to a lower yet stable calibration error and to also approach a converged performance.
Future work. There are several directions we can explore beyond the findings in this paper: (1) External Feedback for Self-Correction. Previous studies show that self-correction with external feedback can improve performance significantly, the difference of it to intrinsic self-correction would be an interesting topic. But acquiring external feedback is expensive particularly if the feedback is from humans, figuring out the performance upper bound of intrinsic self-correction would be helpful for efficiently leverage external feedback. (2) Instruction Optimization. The success of self-correction lies in the injected instruction. Given our findings that the activated concept is the source force driving the convergence of self-correction, it can be used as a supervision signal to search effective instructions. (3) The Connection between In-context Learning and Self-correction. How the in-context learning capability of LLMs helps the emergence of self-correction and how to empower LLMs with a better self-correction capability. (4) The Data-centric Source of Self-Correction. Though previous studies empower LLMs better self-correction capability by learning from self-correction demonstrations [65, 66]. But the most intrinsic source should be from pre-training corpus, which is still unknown.
Reproducibility Statement
This draft aims to reveal how and why intrinsic self-correction can work and enjoys a good property of convergence. We show details of used benchmarks and backbone models, and the prompts are listed in the appendix. Since this draft concentrates on mechanistic analysis, the analysis results can be easily reproduced by following our logics.
Appendix A Additional Experimental Results
Figure 6 shows the results of intrinsic self-correction for the VQA task.